An accelerometer is measuring a structure as described below: Damping ratio of accelerometer = 0.65 Highest frequency measured by the accelerometer = 6000 hz Natural frequency of the accelerometer = 18000 hz What is the percentage error in the accelerometer at the highest frequency it can measure? (Type in a two-decimal number that represents a percentage)
1
Expert's answer
2020-06-10T07:51:10-0400

For finding the error in measurement we can find, "\zeta=0.65" ,"f=6,000,f_n= 18,000"

"E+1 =\frac{1}{(1-\frac{\omega}{\omega_n}^2)^2} + (\frac{2 \zeta \omega }{\omega_n})^2"


"E+1 =\frac{1}{(1-\frac{6000}{18000}^2)^2} + (\frac{2 \times 0.65 \times 6000 }{18000})^2"

E+1 = 1.2656 + 0.1877

E= 0.4533

E= "45.33 %" % this is the error percentage